OpenStudy (anonymous):

If $2543 is invested into a savings account and compounded continuously, what interest rate is required for the investment to triple in 60 years? Round your answer to one decimal place.

OpenStudy (anonymous):

Hello there.

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

Umm...I thought that just meant that we should use A = Pe^rt

OpenStudy (anonymous):

I'm wrong?

OpenStudy (anonymous):

lol Alright...

OpenStudy (anonymous):

So...here's what I tried doing: 7629 = 2543^r60

OpenStudy (anonymous):

Then.... 7629/2543 = e^r60 (sorry, didn't mean to leave out the e)

OpenStudy (anonymous):

Then 3 = e^r60...then ln(3) = ln(e)^r60, then ln(3) = r60

OpenStudy (anonymous):

And presumably ln(3)/60....but apparently, that's wrong

OpenStudy (anonymous):

Any idea what I did wrong?

OpenStudy (anonymous):

i have no idea what i am doing bump this

OpenStudy (anonymous):

Lol Alright then...thanks for your interest, then ;)